# Density of wave fronts

### v1

@@ 7. Existing Literature
The problem of wave fronts is also related to the illumination problem: what is the set of points $Q$ such that the wave front $W_{t}(P)$ never reaches $Q$? This question was first asked the 1950ies. Roger Penrose in 1958, constructed the Penrose unilluminable room. The first polygonal unilluminable room was found in 1995 [32]. In a general polygon, this is not an empty set in general, but it is finite in general [37]. For recent work on translation surfaces see [28]. Note however that having an unilluminable room does not prevent that the wave front can become dense. And having a room that can be illuminated does not prove that fronts become dense. The duck in the center of a circular pond illustrates this.
@@
In Riemannian geometry, a manifold is called a Wiedersehen manifold if for every point $P$ there is a Wiedersehen point $Q$ such that all geodesics through $P$ pass also through $Q$. Wilhelm Blaschke conjectured in 1921 [10], that every Wiedersehen manifold is a sphere. Only in 1963, Green proved that a 2-dimensional Wiedersehen manifold (Wiedersehenfläche) must be isometric to a sphere of constant curvature. See [24] Theorem 5.2.5). Note that for a projective plane of constant curvature, the point $Q$ is the same than $P$ so that it is technically not called a Wiedersehen manifold. A natural question is whether it is true that for any non-Wiedersehen manifold (different from a projective space situation), there exists a point $P$ for which the wave front $W_{t}(P)$ becomes dense and that for any d-manifold with or without boundary that does not allow for a $SO(d-1)$ symmetry, the wave front becomes dense for all points $P$. An example of a $d=2$-manifold with boundary is the unit ball which carries a $SO(1)$ rotational symmetry. An other example is an ellipsoid of revolution which again carries a $SO(1)$ rotational symmetry and for which there exist points $P$ which have a Wiedersehen point $Q$.
@@
We have already mentioned that in general, we expect wave fronts to become dense. What about specific situations in higher dimensions like in the case of billiards in a regular polytop? In the case of a polytop that tesselates Euclidean space we can prove the density by projection. How can one prove for a regular polytop that $W_{t}(P)$ becomes dense? A good place to start would be a solid regular dodecahedron. The proofs given here do not apply. Still, we feel that there should be a simple argument proving that wave fronts are dense in any convex polytop of dimension 2 or larger.
@@
There would be other questions. In the cube case we have seen that the fronts $W_{t}$ becomes disconnected pretty fast. The total length of the curve of course is $2\pi t$. What is the typical length of a connected component, where we only consider connected components $\exp_{t}(I)$ with $I$ an interval in unit sphere $S\subset TM_{P}=\mathbb{R}^{2}$, such that $\exp_{s}(I)$ is a connected curve for all $s\leq t$. These floating pieces get disconnected even more but also start to intersect with other such curves. Is it possible that the set $W_{t}$ can even become connected again for some large $t$?

### v2

@@ 7. Existing Literature
The problem of wave fronts is also related to the illumination problem: what is the set of points $Q$ such that the wave front $W_{t}(P)$ never reaches $Q$? This question was first asked the 1950ies. Roger Penrose in 1958, constructed the Penrose un-illuminable room. The first polygonal unilluminable room was found in 1995 [35]. In a general polygon, this is not an empty set in general, but it is finite in general [40]. For recent work on translation surfaces see [31]. Note however that having an un-illuminable room does not always prevent that the wave front can become dense. (Added: January 26th: The Penrose example already shows that there are regions for which entire parts of the room can not be illuminated. In that case, also the wave front can not be dense. Thanks to Sergei Tabachnikov to point this out to us.) On the other hand, having a room that can be illuminated does not prove that fronts become dense. The duck in the center of a circular pond illustrates this.
@@
In Riemannian geometry, a manifold is called a Wiedersehen manifold if for every point $P$ there is a Wiedersehen point $Q$ such that all geodesics through $P$ pass also through $Q$. Wilhelm Blaschke conjectured in 1921 [11], that every Wiedersehen manifold is a sphere. Only in 1963, Green proved that a 2-dimensional Wiedersehen manifold (Wiedersehenfläche) must be isometric to a sphere of constant curvature. See [26] Theorem 5.2.5. Note that for a projective plane of constant curvature, the point $Q$ is the same than $P$ so that it is technically not called a Wiedersehen manifold. A natural question is whether it is true that for any non-Wiedersehen manifold (different from a projective space situation), there exists a point $P$ for which the wave front $W_{t}(P)$ becomes dense and that for any d-manifold with or without boundary that does not allow for a SO(d1)SO(d-1) symmetry, the wave front becomes dense for all points $P$. An example of a $d=2$-manifold with boundary is the unit ball which carries a $SO(1)$ rotational symmetry. An other example is an ellipsoid of revolution which again carries a $SO(1)$ rotational symmetry and for which there exist points $P$ which have a Wiedersehen point $Q$.
@@
We have already mentioned that in general, we expect wave fronts to become dense. What about specific situations in higher dimensions like in the case of billiards in a regular polytop? In the case of a polytop that tesselates Euclidean space, we can prove the density by projection. How can one prove for a regular polytop that $W_{t}(P)$ becomes dense? A good place to start would be a solid regular dodecahedron. The proofs given here do not apply. Still, we feel that there should be a simple argument proving that wave fronts are dense in any convex polytop of dimension 2 or larger.
@@
There would be other questions. In the cube case we have seen that the fronts $W_{t}$ becomes disconnected pretty fast. The total length of the curve of course is $2\pi t$. What is the typical length of a connected component, where we only consider connected components $\exp_{t}(I)$ with $I$ an interval in unit sphere $S\subset TM_{P}=\mathbb{R}^{2}$, such that $\exp_{s}(I)$ is a connected curve for all $s\leq t$? These floating pieces get disconnected even more but also start to intersect with other such curves. Is it possible that the set $W_{t}$ can become connected again for some large $t$?
Method
7.9.
All illustrations and experiments were done with the computer algebra system Mathematica. Computer code for the illustrations is available online on the Wolfram blog [27].